Shift Summary: 07/03/99 Owl (fwd)

Jason LaDue (jladue@cuhep.Colorado.EDU)
Sat, 3 Jul 1999 06:19:22 -0600 (MDT)

Shift Summary for owl shift, 07/03/99

On Shift: Ugur Erkarslan, Jason LaDue

Hours of beam delivered at all intensities: 4.0h
Hours of beam delivered at nominal intensity: 3.5h
Hours of E832 data (special runs not included): 3.0h
Hours of special runs (laser scans, muon runs...):0.0h

Accounting of lost time:

Various accelerator problems.
Changing tapes, watching runs abort and cleaning tape drives.

========================================================================
Runs ended this shift:
Run 13800 13.0M BEAM_832_NOKTEV1

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift:
None

========================================================================
Overview of shift:
Quiet shift, only one major hang up. No beam to speak of either.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13799 00:21 SEM at 6E12, call the control room and it goes
down to 5E12.

309 00:44 TeV injector synch problems, no beam, est. 20min.

01:01 TeV injector quench, still no beam, est 20min.

339 01:18 Beam returns, trigger soon entires error state.
KFREND fixes by next spill.

363 01:55 Beam goes away.

02:04 Injection Quench at A4 and F2, investigating.

407 02:49 Beam returns for a single spill.

413 02:57 Beam returns, SEM is at 2.1E12 and climbing.

438 03:31 Beam aborts right away and then no beam.

03:37 Control Room calls and says that they are having
problems injecting the beam so the intensity will
be low (~1.6E10) for a while.

04:48 Non-trivial SEM (6E11). It slowly increases.

05:32 Tapes Full and the Run ends.

13800 05:45 I forget to Update the tape numbers causing the
run to abort (DOH!). Not wanting to fresh start
for fear of really screwing things up, I wait
the 10 minutes for the run to abort.

13801 06:02 Run aborts when Drive tps2d1 on plane 3 decides
that now is a good time to be cleaned. I decide
that it is safest to use a new tape instead of
reusing the one that was in there before. The
"old" tape is labeled and put to the side.

13802 06:29 Finally start running smoothly.

13802 21 06:56 CsI channel 1786 has a gain drift of 20%, goes
back down next spill.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift